Given f(x)= 3x-5, g(x)= x2-9. What is (f+g)(-2)


1
Expert's answer
2021-03-31T07:26:06-0400

You have the following functions


"f(x)=3x-5" and "g(x)=x^{2}-9"


The sum in the compound functions is given by


"(f+g)(x)=f(x)+g(x)"


Obtaining the compound function


"(f+g)(x)=f(x)+g(x)\\
(f+g)(x)=3x-5+x^{2}-9\\
(f+g)(x)=x^{2}+3x-5-9\\
(f+g)(x)=x^{2}+3x-14"


Evaluating numerically at x =-2


"(f+g)(-2)=x^{2}+3x-14 \\
(f+g)(-2)=(-2)^{2}+3(-2)-14\\
(f+g)(-2)=4-6-14\\
(f+g)(-2)=-16"


Answer "\displaystyle \color{red}{\boxed{(f+g)(-2)=-16}}"